If a boy completes an assignment in 8 hours and another boy completes the same assignment in 10 hours, then how many hours would it take for both of them to complete the assignment if they work on it together?
1
Expert's answer
2019-10-03T12:36:39-0400

The speed of the first boy:


"\frac{1}{8}"

The speed of the second boy:


"\frac{1}{10}"

The total speed:

"\frac{1}{8}+\frac{1}{10}=\frac{9}{40}"

The new time is


"\frac{1}{\frac{9}{40}}=\frac{40}{9}\ h=4\frac{4}{9}\ h"